OpenStudy (anonymous):

What is the equation of the line that is perpendicular to y = 2x – 4 and has a y-intercept of –2?

OpenStudy (texaschic101):

the slope in the given equation is 2. Do you understand that ?

OpenStudy (anonymous):

yes, so would i find the negative reciprocal?

OpenStudy (texaschic101):

yes...because you are looking for a perpendicular line

OpenStudy (anonymous):

OK! so the answer should be Y= -1/2x+2?

OpenStudy (texaschic101):

not quite...you want the y intercept to be -2

OpenStudy (anonymous):

ok... so y=-2x-2?

OpenStudy (texaschic101):

the y intercept is in the b position of y = mx + b. so your perpendicular equation is y = -1/2 - 2 you got it :)

OpenStudy (anonymous):

ok thanks!

OpenStudy (texaschic101):

YW :) y = -1/2x - 2